OpenStudy (anonymous):

Solve using addition or subtraction method: 5x-9y=-3 4x-3y=6

OpenStudy (anonymous):

i get 9x+12y=3 now what?

OpenStudy (anonymous):

yeah

OpenStudy (anonymous):

multiply second equation bby 3 12x-9y=18 then substract first equation from this , you'll get value of x then substitute back the value of x you'll have y value

OpenStudy (anonymous):

so heres what i get: 5x-9y=-3 12x-9y=18 so it should be -7x=-21?